A women’s shoe department reported the following plans for November: Planned Sales.. $72,000 Planned BOM.. $89,000 Planned Reductions.. $4,000 Planned EOM.. $68,000 Stock on Order (@ Cost).. $13,500 IMU%.. 45% Calculate the buyer’s open to buy (OTB) at cost.
Answer #1

Open to buy (OTB) at cost = OTB value * IMU %

OTB Value = (Planned Sales + Planned EOM + Planned Reductions - Planned BOM) - Stock on Order

=(72,000 + 68,000 + 4,000 - 89,000) - 13,500

= 55,000 - 13,500 = $41,500

Now OTB at cost = $41,500 * 45% = $41,500* 0.45 = $18,675